# Explicit modeling of density dependence in spatial capture-recapture models

### v1

@@ 4 Discussion
In this study, we extended SCR models (e.g., Royle et al.,, 2013, 2018) to estimate the true amount of time of habitat use, an information that is often of interest in ecological research and conservation planning (Hull et al.,, 2016; Dawson et al.,, 2013). In our simulations, we used the robust sampling design to facilitate the inference of habitat use. However, the inference of habitat use does not rely on robust sampling design, which is illustrated by our case study.
@@
Despite the accurate information of habitat use provided by this SCR model, we found it still difficult to estimate the effect of density dependence on apparent survival and recruitment with this model. So far, density dependence is mostly considered at the population level in SCR models (e.g., Chandler and Clark,, 2014); when density dependence is considered at individual level in SCR models, only the locations of territory centroids, but not the information of habitat use, were considered (Milleret et al.,, 2023). The shortcoming of doing so is that individuals will not influence each other through density dependent processes if their territory centroids do not fall at the same site, despite the fact that their territories can still overlap with each other to induce competition. Therefore, we attempted to develop a model that links habitat use information with density dependence. The difficulties in estimating the effect of density dependence in this model is likely due to the fact that the current generation of SCR models can only identify the locations of territory centroids of observed individuals, but not unobserved individuals, under the assumption of uniformly distributed territory centroids (Figure 3). Because the locations of unobserved individuals are misidentified, the model mis-specifies local density, and thus underestimates the effect of density dependence on vital rates. It is likely for the same reason that the model also provides biased estimates of the movement parameter. Overall, our results are in line with the findings of previous studies that showed the limits of SCR models (Sun and Cole Burton,, 2024), and it remains a challenge to link habitat use with vital rates at individual level in SCR models. Further studies are needed to better specify the locations of individuals, probably through the integration of other data resources with SCR data (e.g., Sun et al.,, 2019).
Our study nonetheless showed the importance of separating true habitat use with observation processes. In our case study, we found an unexpected positive density dependence, likely due to missing environmental covariates because higher density is likely to be an indication of better habitat and in turn higher survival rate. Indeed, including environmental covariates to inform habitat use is a key component of this model. Therefore, we highly recommend future studies to consider habitat heterogeneity in SCR models.

### v2

@@ 4 Discussion
In this study, we extended SCR models (e.g., royle2013integrating; royle2018unifying) to estimate the true amount of time of habitat use, an information that is often of interest in ecological research and conservation planning (hull2016habitat; dawson2013habitat). Using both simulations and a case study, we showed that it is possible to separate true habitat use from the observed number of captures under a hierarchical Bayesian Framework. A key component of our model is to allow habitat use be a function of both habitat heterogeneity and distance from territory centroid. Therefore, we highly recommend researchers to obtain and incorporate covariate information to maximize the utility of our model. Another key component of our model is to assume that the total amount of time covered by the survey, $U$, is known, under the assumption of closed population within each primary period. We recommend researchers to carefully select the primary period of their survey to meet this assumption. Further studies could also consider extending the model to allow the violation of this assumption due to, for example, death of an individual during the primary period.
@@
Despite the accurate information of habitat use provided by this SCR model, we found it still difficult to estimate the effect of density dependence on apparent survival and recruitment with this model. So far, density dependence is mostly considered at the population level in SCR models (e.g., chandler2014spatially); when density dependence is considered at individual level in SCR models, only the locations of territory centroids, but not the information of habitat use, were considered (milleret2023estimating). The shortcoming of doing so is that individuals will not influence each other through density dependent processes if their territory centroids do not fall at the same site, despite the fact that their territories can still overlap with each other to induce competition. Therefore, we attempted to develop a model that links habitat use information with density dependence. The difficulties in estimating the effect of density dependence in this model is likely due to the fact that the current SCR model can only identify the locations of territory centroids of observed individuals, but not unobserved individuals without linking the locations of territory centroids to environmental covariates (Figure 3). Because the locations of unobserved individuals are misidentified, the model mis-specifies local density, and thus underestimates the effect of density dependence on vital rates. It is likely for the same reason that the model also provides biased estimates of the movement parameter. In the case study, our model identifies no density dependence on apparent survival and a negative density dependence on recruitment. Because of the model’s tendency to underestimate density dependence, we cannot conclude that there is no density dependence on apparent survival, and studies have suggested that density dependent apparent survival is possible for tigers (bisht2019demography). On the other hand, the negative density dependence on recruitment seems to actually exit (although its strength may be underestimated), which indicates that the tigers may have a lower reproduction and/or immigration rate in areas used by more individuals. For instance, tigers are reported to be reluctant to move into a crowded area to prevent inbreeding depression (smith1991contribution).
Overall, our results are in line with the findings of previous studies that showed the limits of the current generation of SCR models (sun2024importance), and it remains a challenge to link habitat use with vital rates at individual level in SCR models. Therefore, we suggest researcher taking cautions when using this and other SCR models to evaluate density dependence. Further development of SCR models is needed to better identify the locations of individuals, probably through linking them to environmental covariates (e.g., borchers2008spatially) or the integration of other data resources with SCR data (e.g., sun2019incorporating).
Our study nonetheless showed the importance and possibility of separating true habitat use with observation processes in SCR models. Our modeling approach can be used in a broad range of species and populations as long as SCR surveys are applicable. Our model can also be easily adapted to a variety of sampling designs (dupont2021optimal; sun2014trap) or modified to include additional model components such as sex structures (augustine2020sex). Thus, our model is broadly transferable under the framework of spatial statistic trinity.